# Career Modeling with Missing Data and Traces

### v2

@@ 4 General discussion
We have showcased the usefulness of our approach in two applied studies on French civil servants and their trajectory into the private sector, extending the social scientific literature along the way. An innovation in our applied studies is found in the way data were collected and cleaned. Our philosophy was to automate as much work as possible, for example in the name-matching phase, which enabled us to study a population an order of magnitude larger than what is usually done in traditional elite research (prosopography studies rarely exceed a few hundred individuals). However, we were not able to automate everything away, like public-private organization name classification, and the main limitation would again become human time if a researcher attempted to scale up this work by another order of magnitude. Our MCMC sampling procedure is, compared to approximate calculations, more costly, which could also be a problem when trying to scale up this approach. We found that the main cost of the procedure was in the computation of the MLE in the initialization of each Metropolis step, in addition to the random walk which required multiple computations of the likelihood. The cost of the Gibbs step was less important as it was very easy to parallelize it across individuals. Enabling a better scaling of the procedure could enable to study of a larger population, at a finer scale than based on six-month increments.
@@ Acknowledgments
The authors are grateful to four anonymous reviewers, an associate editor, and an editor for their very helpful comments on a previous version of this manuscript. RJR received funding from the European Union under the GA 101071601, through the 2023-2029 ERC Synergy grant OCEAN.
@@ Appendix A Data assemblage
This appendix documents the various procedures we used to assemble the dataset. To assemble data, we queried all traces of the JO linked to the corresponding organizations and manually crafted a rule-based approach to only keep those that define membership. All traces and LinkedIn profiles are attached to a full name. There is therefore a high risk of data being tarnished by homonyms. In this section, we describe the careful procedure used to eliminate this risk as much as possible, while preserving information.
A.1 Selection of administrative traces
$\frac{\mathbb{P}[O|S_{1}]}{\mathbb{P}[O]}=\frac{\#\{\text{traces linked to org}\ O\}\cap S_{1}}{\sum_{x}\#\{\text{traces linked to org}\ x\}\cap S_{1}}/\frac{\#\{\text{traces linked to org}\ O\}}{\sum_{x}\#\{\text{traces linked toorg}\ x\}}$
which can be interpreted as a relative risk score. We only keep traces linked to organizations with a score above 2, or above 1 if they have fewer than 1000 traces, or above 0.5 if they have fewer than 100 traces. This effectively removes organizations such as the army or the police where many promotions are described for the entire force, and constitute the main sources of perfect homonyms in $S_{1}$. For traces that are not linked to any organization, we also include them if they are a “délégation de signature” (meaning an authority delegates administrative power to the civil servant, a practice that remains limited in scope) or if it states an admission into a “grand corps”. This enables a good tradeoff between minimizing the homonymy risk and keeping as much data as possible, and yields a second set of traces $S_{2}$. For the first population, we then manually search for every individual for which we cannot find any trace upon graduation, and document the cases of name-changing for individuals who got married during their time as students (around 200 individuals) and add the relevant traces to the total after filtering them using the same procedure. For the second population, if a trace in $S_{2}$ documents the person has had another name, we also add the relevant traces after passing them through the same filters as before (this also represented around 200 individuals). This yields our final set of traces $S_{3}$.
A.2 Linkage with digital profiles
For each person in our population, we search on LinkedIn for up to three profiles that share the same name, download them, and then only keep the ones that list an experience in our starting set of organizations. For example, if an individual is included because they are part of the Statistical office, we only keep profiles with such an affiliation, using a rule-based approach. When more than one profile qualifies, we score the candidate profiles based on their affiliations and only keep the one with the highest score. Overall, we designed our scheme to minimize false matchings between sources rather than maximizing the number of profiles that would be matched; manual inspection shows that it is very robust to homonymy problems in the matching between the two sources. The proportion of individuals that could be matched with a digital profile for subgroups can be found in Appendix B.
For each profile linked to an individual, we obtain a list of dated professional positions that are shared between the public and private sectors (or “spheres”, see Daho and Gally (2018); France and Vauchez (2017)). We classify organization names between public and private positions, using a rule-based method, progressively adding filters to capture all public organizations. This strategy enabled us to cover manually the entire sample. Companies entirely owned by the state (e.g. La Poste, France Television) are also coded as public, alongside companies with the “EPIC” status (e.g. SNCF, OSEO). Contrary to Daho and Gally (2018), companies that only have the state as a minority shareholder (e.g. Renault, TotalEnergies) are coded as private. In addition, positions at Universities and research bodies are not coded as either public or private: it appears they correspond to associate teaching positions that do not constitute the main occupation for people in our population, and so we leave them out. This reasoning leads us to also not code implication in a political campaign, as this is often superposed to other engagements. The same goes for positions in associations, naturally including charities, but also positions in unions, which include corporate representation bodies such as Medef. Finally, to ensure homogeneity with Journal Officiel, we do not include positions in world organizations (e.g. World Health Organization). This strategy of not coding uncertain cases is enabled by our modeling approach, as it doesn’t require full trajectories, and reduces the impact of such decisions on results. Without claiming this coding scheme is perfect, these decisions only have a minor impact on the substantial research findings, and all 1488 coding filters will be made fully accessible in the appendix for re-use.
A.3 Exact perimeter for organizational belonging
Readers interested in the functioning of the French “Grands Corps” will probably be interested in our precise identification of traces that define membership in any of them. Our definition is built in an empirical manner even if there are two defining principles. On the one hand, the definition is heavily constrained by the nature of the archive and what it documents, and on the other hand, it is based on a continuous discussion with sociologists and political scientists so the criterion adopted here would match usual accounts, and avoid being overly broad or narrow.
Concretely, we begin by surveying the Steinertriples API for the different groups we are interested in. We first filter out traces depending on the organization and their nature, using a rule-based approach summarized in the table (0 meaning we filter out). This was defined by discussing with field experts with concrete examples of the different cases.
@@ A.3 Exact perimeter for organizational belonging
  • For ce, traces that are not affectation, we only keep traces matching (auditeur)| (auditrice)|(conseiller)|(conseillère)|(maître). We focus on Conseillers d’États and auditors.
@@
Appendix B Descriptive measures
@@
Table 6: Number of individuals affiliated by service (1990-2022).
@@
Table 7: Proportion of individuals by group with a matched profile.
@@
Appendix C Additional robustness checks
C.1 Choice of $\varphi$
In our implementation, we include the covariates $A(1)$ and $A(\varphi)$ with $\varphi=0.8$. We performed additional tests with different weighting parameters $\varphi$ for the main model of interest (study 1, model 3) $\varphi=0.6,0.7,0.8,0.9$, and compared with the reconstruction loss criterion ($0.222,0.239,0.216,0.229$). The best-performing model , is when choosing $\varphi=0.8$.
@@ C.1 Choice of $\varphi$
C.2 Data heterogeneity concerns and modular inference
In our inference procedure, we use both values of $(X_{t}^{i})$ which have been observed (on LinkedIn), and values which have been simulated from the appropriate conditional; both types of values will be used to infer the parameters of interest. As our two data sources are of very different nature, the heterogeneity it may create can be disquieting. Even in the slightest misspecification case, simulated data could be a lot less trustworthy than observed data, and errors on the former could be propagated to the entire estimation procedure. In addition, even if we attempted to classify digital data so it would match as best as possible administrative accounts, some residual artefactual differences could still be observed.
This problem of having multiple, heterogeneous data modules is common and has been discussed in the Bayesian literature Jacob et al. (2017). For each configuration, we consider two settings: (a) all values are treated equally; (b) simulated values are given a weight of 0.01, whereas observed values are given a weight of 1. This second strategy effectively grants all importance to individuals who have a digital profile.
Simulations in modular inference control condition leads to a similar conclusion on $\gamma^{1}$, with a posterior on $\gamma^{0}$ a little more shifted to the negative side.
Appendix D Digital appendix and data availibility
The scripts and original data are available online at https://github.com/theovoldoire/Hidden_Careers_Public. The MCMC output for all models mentioned in the paper is available online at https://zenodo.org/doi/10.5281/zenodo.10365276.

### v3

@@ 6 Discussion
Using this method, we conclude that exit probabilities from the state have not increased on the period 1990-2022 for ENA graduates, contrary to what we expected before conducting this study. The data rather show that the exit probability has remained constant or has possibly decreased slightly. Furthermore, the probability for a former civil servants to come back to work in the public sector after leaving has significantly increased over the period. In other words, when civil servants leave for the private sector, they now have a higher probability of coming back. This is notably a contribution with respect to (bouzidi_pantouflage_2010), who stressed that they could not study such evolutions given the constraints of their data despite the demands of their journal referees, and showcases the advantages of studying larger groups via data augmentation. In addition, if there seems to be an interaction between political majority changes with election years and revolving doors for this population, it seems to affect the probability of coming back from the private sector than the probability of leaving.
Across the whole elite French civil service, we find three profiles of organizations depending on the probability of leaving and the probability of coming back: one with high probability of leaving but low probability of coming back; one with low probability of leaving but high probability of coming back; and Inspection of Finances standing out from the rest with high probability of both leaving and coming back. Women have overall lower probability of leaving for the private sector than men, but the situation is not constant across public organizations.
Our approach has allowed us to extend the social science literature on French civil servants and their trajectories into the private sector. An innovation in our applied studies is found in the way data were collected and cleaned. Our philosophy was to automate as much work as possible, for example in the name-matching phase, which enabled us to study a population an order of magnitude larger than what is usually done in traditional elite research (prosopography studies rarely exceed a few hundred individuals). However, we were not able to automate everything away, like public-private organization name classification, and the main limitation would again become human time if a researcher attempted to scale up this work by another order of magnitude. Our MCMC sampling procedure is more costly than approximate calculations: calculations remained very manageable in our study, but could also become an issue when trying to scale up this approach. Enabling a better scaling of the procedure could enable to study of a larger population, and at a finer scale than based on six-month increments.
Data and code availability
The scripts and original data are available online at https://github.com/robinryder/career-paths-public. The MCMC output for all models mentioned in the paper is available online at https://zenodo.org/doi/10.5281/zenodo.10365276.
@@ Acknowledgments
The authors are grateful to six anonymous reviewers, an associate editor, and an editor for their very helpful comments on previous versions of this manuscript. RJR received funding from the European Union under the GA 101071601, through the 2023-2029 ERC Synergy grant OCEAN.
@@ Appendix A Data assemblage
This appendix documents the various procedures used to construct the population and datasets. Subsection A.1 describes how the population is defined based on a thoroughly constructed subset of administrative traces in JORF. Subsection A.2 describes how, more generally, decisions in JORF are handled to construct the trace data source. This process is distinct from the population definition task done in subsection A.1 and relies on a much wider array of administrative traces. Subsection A.3 describes how linkage to the detailed data source (Linkedin) is performed. The procedure implemented to handle the risk for homonyms and false positive matchings is described in detail.
A.1 Exact perimeter for organizational belonging
To construct the population in this article, a restrictive set of traces from JORF is used, with much more labor done by hand and back and forth in discussion with sociologists. This enabled to arrive at a reasonably grounded definition of what constitutes membership in any of the groups discussed in the article. Given the complexity of the legal organization of French Grands corps, it is reasonable that other definitions could be reached or argued for, and our choice is thus to exhaustively share here the decision-rule system that was used to define membership in a group.
We expect that a social scientist working on any specific group could motivate a finer definition of organizational membership, with further distinctions to be made within groups. Readers should thus keep in mind that our substantive contribution is essentially comparative, and that further variation ought to be discovered at a finer level. Further, readers should also note that the definitions are necessarily constrained by the nature of the archive and what it documents. If a career transition cannot legally be found in the JORF, then, it cannot be used to define populations in the manner proposed here. This is a problem for many trajectory studies, but since for elite civil servants, nomination procedures are on the highly-formalized end of the spectrum, we find that problem to be manageable. However, it is clear that the following definitions are, like the organizational boundaries they are trying to approximate, socially constructed and context-dependent. We find there is already a contribution in making the decision rule transparent so that it can be further improved or specialized for other studies.
Concretely, we begin by surveying the Steinertriples API for the different groups selected for the study. Traces are first filtered depending on their tagged organization and their nature, using a rule-based approach summarized in Table 4 (0 means filtered out, 1 means kept in). This was guided using substantive knowledge of civil servants and asking some help from sociologists. Some decisions were taken by studying specific biographies of individuals that would be included or excluded from the lists by any decision, to see if it would correspond to an intuitive definition of membership.
@@ A.1 Exact perimeter for organizational belonging
  • For ce, traces that are not affectation, we only keep traces matching (auditeur)| (auditrice)|(conseiller)|(conseillère)|(maître). We focus on Conseillers d’État and auditors.
@@
A.2 Selection of administrative traces
[…]
which can be interpreted as a relative risk score. We only keep traces linked to organizations with a score above 2, or above 1 if they have fewer than 1000 traces, or above 0.5 if they have fewer than 100 traces. This effectively removes organizations such as the army or the police where many promotions are described for the entire force, and constitute the main sources of perfect homonyms in $S_{1}$. For traces that are not linked to any organization, we also include them if they are a “délégation de signature” (meaning an authority delegates administrative power to the civil servant, a practice that remains limited in scope) or if it states an admission into a “grand corps”. This enables a good tradeoff between minimizing the homonymy risk and keeping as much data as possible, and yields a second set of traces $S_{2}$. For the first population, we then manually search for every individual for which we cannot find any trace upon graduation, and document the cases of name-changing for individuals who got married during their time as students (around 200 individuals) and add the relevant traces to the total after filtering them using the same procedure. For the second population, if a trace in $S_{2}$ documents the person has had another name, we also add the relevant traces after passing them through the same filters as before (this also represented around 200 individuals, which were also all checked by hand). This yields our final set of traces $S_{3}$. This set is the one used as the “traces” dataset throughout the article.
A.3 Linkage with digital profiles
For each profile linked to an individual, we obtain a list of dated professional positions that are shared between the public and private sectors (or “spheres”, see daho_cabinet_2018; france_sphere_2017). We classify organization names between public and private positions, using a rule-based method, progressively adding filters to capture all public organizations. This strategy enabled us to cover manually the entire sample. Companies entirely owned by the state (e.g. La Poste, France Télévisions) are also coded as public, alongside companies with the “EPIC” status (e.g. SNCF, OSEO). Contrary to daho_cabinet_2018, companies that only have the state as a minority shareholder (e.g. Renault, TotalEnergies) are coded as private. In addition, positions at Universities and research bodies are not coded as either public or private: it appears they correspond to associate teaching positions that do not constitute the main occupation for people in our population, and so we leave them out. This reasoning leads us to also not code implication in a political campaign, as this is often superposed to other engagements. The same goes for positions in associations, naturally including charities, but also positions in unions, which include corporate representation bodies such as Medef. Finally, to ensure homogeneity with Journal Officiel, we do not include positions in world organizations (e.g. World Health Organization). This strategy of not coding uncertain cases is enabled by our modeling approach, as it doesn’t require full trajectories, and reduces the impact of such decisions on results. Without claiming this coding scheme is perfect, these decisions only have a minor impact on the substantial research findings, and all 1488 coding filters will be made fully accessible in the appendix for re-use.
Appendix B Further descriptive statistics
The following tables provide descriptive statistics. Table 5 describes the population size of ENA students included each year in the study. Table 6 describes the population size of the different groups of elite civil servants (note that some individuals may be counted twice). Table 7 gives the proportion of individuals, in any of the groups that are matched to a digital profiles. Table 8 describes the gender proportion and the proportion of ENA graduates by organization.
@@ Appendix B Further descriptive statistics
Table 6: Number of individuals affiliated by group (1990-2022).
@@
Table 7: Proportion of individuals by group with a matched profile on LinkedIn.
@@
Appendix C Additional robustness checks in the choice of $\varphi$
In our implementation, we include the covariates $A(1)$ and $A(\varphi)$ with $\varphi=0.8$. We performed additional tests with different weighting parameters $\varphi$ for the main model of interest (model 3 in Section 5.2.1) $\varphi=0.6,0.7,0.8,0.9$, and compared with the reconstruction loss criterion ($0.222,0.239,0.216,0.229$). The best-performing model is when choosing $\varphi=0.8$.
@@ Appendix C Additional robustness checks in the choice of $\varphi$
Appendix D Digital appendix and data availability
The scripts and original data are available online at https://github.com/robinryder/career-paths-public. The MCMC output for all models mentioned in the paper is available online at https://zenodo.org/doi/10.5281/zenodo.10365276.
A note on the type of data released with this paper is in order. The released data is at an individual level with a 6-months precision frame, and is fully anonymized. Only abstract numeric identifiers are released, alongside the following information: gender, public organization group covered in the study (e.g. ENA, Conseil d’État), whether a trace in JORF was observed in a given 6-months time frame, whether the identifier is linked to a digital profile, whether the person is working in the public or private sector in a given six-months time frame according to their digital profile. All other columns of the dataframe released are computed from this basic information. This is exactly the strictly required information for the models trained and presented in the paper, and release is mandatory so that the proposed statistical techniques are transparent and replicable.
To avoid risks of de-anonymization, no other information is released or can be released to the public. For example, specific private organization names or public organization names besides the exact list studied in the paper are not released. This, alongside the time granularity of six months, ensures that the cost of de-anonymization of this study’s data is lower than the cost of assembling the data from the ground up, making de-anonymization an essentially useless endeavor. The information, despite being linked to an individual identifier, is also sufficiently crude so that matching it to another dataset cannot provide any substantial advantage, making de-anonymization more than just costly, but also without benefit. For this reason, it is also not possible to share the scripts that were used to originally collect the data, especially for the digital profiles part, as sharing them would be essentially equivalent to releasing the uncensored data. Even if the whole data pipeline run at any given time relies exclusively on publicly available information, it still takes as input individual data. As such, only the statistical inference scripts are made public and no other scripts can be made public.
Finally, despite it being very clear from the paper’s framing, let us explicitly state that the statistical analysis is purely motivated from a structural, scientific, analysis of public-private paths of elite civil servants. Data was collected without any interest for any specific individual, and the sole interest was to describe trends. It is anchored in a strong statistical and social scientific framing, and is also motivated by a clear public interest question, that of the deontology of elite civil servants and the breadth of revolving doors in France put in the historical context of a 30 years time frame.